# Analyzing Classroom Interaction Data Using Prompt Engineering and Network Analysis

### v1

@@ 1.1 Background: Classroom Dialogue Analysis
Classroom interactions play a crucial role in the learning process and substantially influence students’ motivation, comprehension, and academic achievement. Through effective student-teacher and student-student interactions, learners develop essential skills in critical thinking, collaborative work, and problem-solving. Analysis of these student interactions reveals underlying behavioral patterns that either facilitate or impede effective learning.
@@
While classroom conversations offer valuable insights into learning dynamics, their analysis presents considerable analytical challenges due to their inherent complexity and high dimensionality. These conversations are often characterized by dynamic shifts in topic, speaker turns, and the interplay of various communicative functions, making manual analysis particularly difficult. These also encompass speaker characteristics (e.g., their speaking style, experience), emotional tone (e.g., enthusiasm, uncertainty, affect), and context (e.g., type of task, classroom environment). To address such complexity, researchers often simplify the analysis by categorizing the intent of individual messages or utterances using predetermined coding schemes. For example, researchers may classify utterances according to function, such as whether they are questions, answers, elaborations, or agreements (Webb et al., 2014; Franke et al., 2015). This strategy entails classifying utterances based on semantic, syntactic, and pragmatic features, and using these classifications to gain insights into specific aspects of the interaction.
@@
Strategic prompt engineering enables researchers to classify dialogue interaction types even without traditional training data for fine-tuning, provided the prompts incorporate sufficient contextual information. Nevertheless, while prompt engineering facilitates classification of individual utterances, it cannot fully capture the complex dynamics of classroom interactions. The classification of discrete utterances reveals speaker identities and discussion content but fails to illuminate critical patterns such as interaction frequency, structural relationships, and the relationships between students.
@@ 1.2 Purpose and Structure of the Paper
In this paper, we present an integrated framework for analyzing classroom interactions by combining the strengths of prompt engineering and network analysis. Our methodology begins with automated dialogue classification through prompt engineering, addressing the limitations of traditional manual labeling approaches. We employ carefully constructed prompts to guide the generative model in classifying interactions according to predefined criteria, establishing the foundation for network-based analysis of classroom dynamics. We then transform these classified interactions into network representations and analyze them using a network mediation model. Specifically, we employ the additive and multiplicative effects network (AMEN) model (Hoff, 2021) model to extract latent social structures and utilize them as mediators to investigate how classroom networks mediate the relationship between gender and mathematics performance. Through examination of how these latent social structures influence student outcomes, with particular attention to gender differences in mathematics performance, we illuminate the mechanisms by which network structures mediate the impact of classroom dynamics on student achievement. This comprehensive methodological integration can provide educational researchers with a powerful analytical tool for investigating the complex relationships between social dynamics and academic performance.
@@ 2.1 Data: Classroom Interaction Dataset
Student C: Mine got 35.
@@ 2.2 Prompt Engineering Methodology
The complexity and high-dimensional nature of classroom interactions necessitate a robust method for classifying and analyzing dialogue intent. As discussed in the Introduction Section, existing analytical approaches require manual labeling, which is resource-intensive, while deep learning methods require extensive labeled datasets. Our methodology addresses these issues through prompt engineering, by leveraging pre-trained models to classify dialogue intents accurately without labeled training data or computational fine-tuning. This approach maximizes the capabilities of pre-trained models while maintaining efficiency in resource utilization and processing time.
@@
In-context learning (ICL) is a specialized application of prompt engineering in which pre-trained large language models (LLMs) learn exclusively from information provided within the prompt, eliminating the need for parameter updates. This approach parallels human learning processes through inference while minimizing computational requirements for task adaptation (Dong et al., 2022). Few-shot prompting, one of the ICL strategies, enhances model comprehension by incorporating targeted task examples within the prompt. The effectiveness of few-shot prompts in ICL has been explained through probabilistic frameworks such as hidden Markov models (Baum and Petrie, 1966), where carefully selected examples guide model reasoning by establishing sequential dependencies that reduce prediction uncertainty (Xie et al., 2021). Min et al. (2022) further demonstrates that optimal performance depends not merely on increasing example quantity but on developing prompt formats that precisely align with task-specific requirements, including both structural elements and linguistic formulation.
@@
Figure 1: System and User Prompts for Utterance Classification. Includes guidelines, category definitions, and example dialogues for analyzing and classifying the contribution of utterances in collaborative problem-solving processes.
Figure 1 illustrates the user and system prompts implemented in this study. The prompts define the utterance classification task, outline categorization rules, and provide illustrative dialogues to clarify distinctions between categories. The user prompt provides clear guidelines for classifying utterances into three main categories: ‘explain own ideas’, ‘engage others’ ideas’, or ‘uncorrelated’. For the ‘engage others’ ideas’ category, it further defines three levels of engagement (low, medium, high) and offers specific criteria for making classification decisions, including special cases and exceptions. The system prompt supports these guidelines by stressing the importance of accurate categorization, emphasizing subtle differences between labels, and ensuring consistency across classifications. To enhance output reliability and reduce variability in the model’s responses, we adjusted the temperature parameter to nearly zero, optimizing consistency in classification.
2.3 Utterance Classification Results
We conducted utterance classification using prompts carefully constructed, as illustrated in Figure 1. These prompts were designed incorporating comprehensive contextual information, including dialogue tasks, categorization rules, and classification guidelines. To optimize classification accuracy, we included five few-shot examples in the prompts, each structured to specify speaker identity, utterance content, classification label, engagement level, and relevant references. Complete prompt specifications, including the few-shot examples, are detailed in the Section C of the Supplementary Material. Instead of individual utterances, dialogue chunks segmented by timestamps from the original dataset were included as the inputs for classification within our prompts. This approach allows the prompts to capture conversational context effectively and inform the classification decisions. The classification process was executed on an NVIDIA RTX A6000 GPU with an Intel(R) Xeon(R) Gold 6426Y CPU in a Linux environment, requiring approximately 1.5 hours for completion.
[…]
  Utterance Type Engagement Level Number of Utterances
Explain Own Idea (EXP) 1,335
Engage Others Idea (EOI) Total 1,251
High Engagement 107
Medium Engagement 553
Low Engagement 591
Uncorrelated 1,544
Total Utterances 4,130
[…]
  Name texts Label Engagement Level
  Delilah Okay now. First I put 2 days to represent the days. And then I put 70 units in each group (gestures across the two groups on her notebook). explain own idea -
Keila First I… uncorrelated -
Delilah And I came up with 39 because each one had 39. explain own idea -
Keila Did you put it one by one? engage others idea medium
Delilah Yeah. explain own idea -
Keila Mine got 35. engage others idea low
Delilah So I got 1, 2, 3, 4, 5, 6, 7, 8, blah blah blah. (counting the marks in one group) explain own idea -
Keila Did you put 1, 2, 3, 4, 5, 6? (gestures marking in alternating groups as she counts, demonstrating action in solving) engage others idea medium
Delilah Yeah! And 10 divided by 2 is 3. I meant 39. explain own idea -
Table 2: Examples: Identifying labels (explain vs. engage) and engagement levels.
  Name texts Label Engagement Level
@@ 2.3 Utterance Classification Results
William Wait hold on. You don’t add 7 plus 7. If you add 7 plus 7 that would be 14. But are you trying to add? Or times? engage others idea high
@@
Second, Table 3 illustrates the system’s ability to capture complex interaction sequences. The example begins with a student’s explanation of their approach (classified as EXP), followed by a peer’s substantive disagreement (classified as high-level EOI). The subsequent probing question, which builds upon this disagreement, maintains the high-level EOI classification. This sequence demonstrates the capacity of the system to identify sophisticated interactions where students progressively build upon and critically engage with each other’s mathematical reasoning.
  Name texts Label Engagement Level
@@
Natalie I’m looking at the hundreds chart. (refers to yellow sheet) And then I went to 75. explain own idea -
@@
Kimberly It was 70. engage others idea low
Natalie On the second strategy. explain own idea -
@@
Kimberly We are at the strategies, but she and me realized that we strategies on the 200 chart. uncorrelated -

### v2

@@ 1.1 Background: Classroom Dialogue Analysis
Classroom interactions play a crucial role in the learning process and substantially influence student motivation, comprehension, and academic achievement. Through effective student-teacher and student-student interactions, learners develop essential skills in critical thinking, collaborative work, and problem solving. Analysis of these student interactions reveals underlying behavioral patterns that facilitate or impede effective learning.
@@
While classroom conversations offer valuable insights into learning dynamics, their analysis presents considerable analytical challenges due to their inherent complexity and high dimensionality. These conversations are often characterized by dynamic shifts in topic, speaker turns, and the interplay of various communicative functions, making manual analysis particularly difficult. They also encompass speaker characteristics (e.g., their speaking style, experience), emotional tone (e.g., enthusiasm, uncertainty, affect), and context (e.g., type of task, classroom environment). To address such complexity, researchers often simplify the analysis by categorizing the intent of individual messages or utterances using predetermined coding schemes. For example, researchers may classify utterances according to function, such as whether they are questions, answers, elaborations, or agreements (Webb et al., 2014; Franke et al., 2015). This strategy entails classifying utterances based on semantic, syntactic, and pragmatic features, and using these classifications to gain insights into specific aspects of the interaction.
@@
Building on these capabilities, recent studies have explored prompt engineering for automating classroom dialogue analysis. Long et al. (2024) reported that LLMs like GPT-4 can perform coding tasks up to 30 times faster than human researchers while achieving over 90% agreement for relatively straightforward classification tasks. On the other hand, Ganesh et al. (2024) found that while models perform well on tasks with clear linguistic cues, they exhibit unstable and inconsistent performance on complex, theoretically motivated qualitative coding tasks requiring nuanced interpretation. Ganesh et al. (2024) added that fine-tuned deep learning models (i.e., models with parameters updated using labeled training data) often outperform prompt-based approaches, which rely solely on prompt design without parameter updates. These findings highlight the need for methodological innovations that can enhance prompt-based classification reliability while preserving the flexibility advantages of not requiring extensive pre-labeled datasets.
Despite these challenges, strategic prompt engineering still enables researchers to classify dialogue interaction types without traditional training data for fine-tuning, provided the prompts incorporate sufficient contextual information. Nevertheless, even with enhanced prompt engineering approaches, a fundamental limitation remains: while prompt engineering facilitates classification of individual utterances, it cannot fully capture the complex relational dynamics of classroom interactions. The classification of discrete utterances reveals speaker identities and discussion content but fails to illuminate critical patterns such as interaction frequency, structural relationships, and relationships among students.
@@ 1.2 Purpose and Structure of the Paper
In this paper, we present an integrated framework for analyzing classroom interactions by combining the strengths of prompt engineering and network analysis. Our methodology begins with automated utterance classification through prompt engineering, addressing the limitations of traditional manual labeling approaches. We employ carefully constructed prompts to guide the generative model in classifying interactions according to predefined criteria, establishing the foundation for network-based analysis of classroom dynamics.
To enhance classification reliability, we employ an ensemble strategy that aggregates the predictions of multiple instruction-tuned LLMs using majority voting. This simple yet effective method produces more consistent labels (Yue et al., 2024) and helps mitigate limitations of individual models, such as biases, sensitivity to prompts, and task-specific variability (Jiang et al., 2023; Lu et al., 2024; Chen et al., 2025). By building consensus across models, the ensemble approach enables more robust utterance classification and improves the reliability of subsequent network analyses compared to relying on a single model.
Further, we incorporate an uncertainty quantification framework to systematically detect ambiguous utterances that may benefit from human review. This involves assessing disagreement among models and evaluating token-level probability distributions to gauge labeling uncertainty. Utterances with high entropy or inconsistent probability patterns across models are flagged for expert inspection. In this way, annotation efforts can be allocated more efficiently while ensuring strong reliability and accuracy in classification.
We then transform these classified interactions into network representations and analyze them using a network mediation model. Specifically, we employ the additive and multiplicative effects network model (AMEN) (Hoff, 2021) to extract latent social structures and use them as mediators to investigate how classroom networks mediate the relationship between gender and mathematics performance. Through examination of how these latent social structures influence student outcomes, with particular attention to gender differences in mathematics performance, we illuminate the mechanisms by which network structures mediate the impact of classroom dynamics on student achievement. This comprehensive methodological integration can provide educational researchers with a powerful analytical tool to investigate the complex relationships between social dynamics and academic performance.
@@ 2.1 Data: Classroom Interaction Dataset
Student C: Mine got 35. I think you missed one group.
@@ 2.2 Prompt Engineering Methodology
The complexity and high-dimensional nature of classroom interactions require a robust method to classify and analyze the intent of dialogue. As discussed in Section 1, existing analytical approaches require manual labeling, which is resource-intensive, while deep learning methods require extensive labeled datasets. Our methodology addresses these issues through prompt engineering, by leveraging pre-trained models to classify dialogue intents accurately without labeled training data or computational fine-tuning. This approach maximizes the capabilities of pre-trained models while maintaining efficiency in resource utilization and processing time.
@@
In-context learning (ICL) is a specialized application of prompt engineering in which pre-trained LLMs learn exclusively from information provided within the prompt, eliminating the need for parameter updates. This approach parallels human learning processes through inference while minimizing computational requirements for task adaptation (Dong et al., 2024). Few-shot prompting, one of the ICL strategies, enhances model comprehension by incorporating targeted task examples within the prompt. The effectiveness of a few-shot prompts in ICL has been explained through probabilistic frameworks such as hidden Markov models (Baum and Petrie, 1966), where carefully selected examples guide model reasoning by establishing sequential dependencies that reduce prediction uncertainty (Xie et al., 2021). Min et al. (2022) further demonstrates that optimal performance depends not merely on increasing the amount of examples, but on developing prompt formats that precisely align with the specific requirements of the task, including both structural elements and linguistic formulation.
In our study, the prompts are systematically designed using a two-step chain-of-thought structure to improve classification accuracy and consistency. The first step generates explicit reasoning for the classification decision, while the second step applies this reasoning to produce the actual categorization. The system prompt incorporates comprehensive classification criteria and categorization rules that define the utterance classification task. Specifically, it provides explicit guidelines for categorizing interactions as “explaining own ideas” or “engaging with others’ ideas,” with the latter further differentiated into low, medium, or high engagement levels. In addition, the system prompt outlines specific criteria for making classification decisions, including handling special cases and exceptions. Throughout prompt design, we emphasize the importance of accurate categorization and consistency across classifications to ensure that subtle distinctions between labels are properly recognized and maintained.
The user prompt complements this framework by providing essential contextual information required for classification decisions, including previous utterances and speaker information that enable the model to understand the flow of conversation and the relationships between participants. This contextual grounding facilitates more nuanced classification of utterances into the three main categories: explain your own ideas, engage with others’ ideas, or uncorrelated. The prompts incorporate illustrative dialogues that clarify distinctions between categories and demonstrate practical application of classification rules. Detailed prompts are provided in Section LABEL:supp-sec:prompt of the Supplementary Material. To enhance output reliability and reduce variability in model responses, we configure the temperature parameter to zero, ensuring reproducible results for our analysis.
To enhance classification robustness and reduce individual model biases, we employ an ensemble of seven instruction-tuned LLMs comprising both open-source models (Meta’s Llama-3.1-8B (Dubey et al., 2024), Alibaba’s Qwen-3-32B-AWQ (Yang et al., 2025), and Microsoft’s Phi-4 (Abdin et al., 2024)) and commercial models (OpenAI’s GPT-4.1 (OpenAI, 2025), Google’s Gemini-2.5-Pro (Google DeepMind, 2025a), Gemini-2.5-Flash (Google DeepMind, 2025b), and Anthropic’s Claude-4-Sonnet (Anthropic, 2025)). These models exhibit distinct response patterns due to variations in training data, fine-tuning approaches such as reinforcement learning from human feedback and supervised tuning, and underlying development methodologies. Rather than relying on predictions from a single model, we collect responses from all ensemble members and assign the most frequently occurring classification as the final label through majority voting. This ensemble approach mitigates individual model biases and produces more robust classification labels through cross-model consensus.
All models are prompted using a standardized few-shot format that incorporates task-specific interaction examples designed to improve instruction adherence and classification accuracy. Prompt construction adheres to the official formatting guidelines for each model to ensure methodological consistency and reproducibility across the entire ensemble. Detailed model descriptions and formatting guidelines are provided in Section LABEL:supp-sec:llm of the Supplementary Material. This standardized prompting approach ensures equitable comparison and consistent behavioral patterns across all ensemble members, effectively mitigating the influence of model-specific idiosyncrasies on classification outcomes.
2.3 Utterance Classification Results
The classification process implements a two-step chain-of-thought structure, requiring models to first generate explicit reasoning before producing the final categorization. Each few-shot example demonstrates comprehensive reasoning processes covering speaker identification, utterance content analysis, classification labeling, engagement level assessment, and relevant contextual references. Complete prompt specifications, including few-shot examples, are provided in Section LABEL:supp-sec:prompt of the Supplementary Material. The classification process was executed on an NVIDIA RTX A6000 GPU with an Intel Xeon Gold 6426Y CPU in a Linux environment.
Following the utterance classification, we excluded entries labeled as “uncorrelated” from subsequent analyses. The remaining utterances, classified as EXP and EOI, were segregated to construct distinct interaction networks for each type of discourse. As a result, a total of 4,233 utterances were analyzed, 497 classified as EXP and 696 as EOI. The latter category was further divided into three subcategories based on engagement level: high, medium, and low. Table 1 presents a detailed breakdown of these classifications.
  Utterance Type Engagement Level Number of Utterances
Explain Own Idea (EXP) 497
Engage Others Idea (EOI) Total 696
High Engagement 103
Medium Engagement 205
Low Engagement 388
Uncorrelated 3,040
Total Utterances 4,233
  Name Utterances Label Engagement Level
  Kimberly I thought you did just it by 5 because like (points to N’s notebook). engage others idea low
Natalie Then I got to 5s and made 10, so I put 5 and 5 to (inaudible). And then I counted how much there was both and…it was 35. explain own idea -
Kimberly Okay. Samantha go next. uncorrelated -
Samantha What I did was right here. I did yesterday and then I didn’t have enough for everyone. And then I passed the last 10 because if I put it in with these, wouldn’t it be equal and then I split it. And then I carried it in and then I divided by once, and then I counted it and then I got 1, 2, 3, 4, 5 and this was 30, and then I got 35. explain own idea -
Kimberly What I did was like I did like Samantha’s but I just labeled one day and two day, and then I just made a cross. Then, here I did the same thing, I got the seven 10s, but when I end up with 2, I just cross 2 off and then the 10. But then I noticed this was like that so I turned them into little 10 ones. And then I got 1, 2, 3, 4, 5, 6, 7, 8, 9. 10. I got 35. explain own idea -
Natalie (has a look of disbelief) How do you know it was 35? engage others idea medium
Kimberly Like I counted it. 10, 20. I already knew this was 20, 30 explain own idea -
Natalie 40. uncorrelated -
Kimberly (laughs)…30, 31, 32, 33, 34, 35. And that’s my answer for how I got to 35. (adjusts camera) And now can I go first for the second number share? explain own idea -
Natalie Yes. uncorrelated -
Table 2: Examples: Identifying labels (explain vs. engage).
@@ 2.3 Utterance Classification Results
  Name Utterances Label Engagement Level
@@
William Wait hold on. You don’t add 7 plus 7. If you add 7 plus 7 that would be 14. But are you trying to add? Or times? engage others idea medium
@@
Table 3 illustrates the ability of the system to capture complex interaction sequences. The example begins with a student’s explanation of their approach, which is classified as explain own idea. This is followed by a peer’s substantive disagreement, appropriately classified as engage with others’ ideas with high engagement level. The subsequent question of probing, which builds on this disagreement, is classified as Engage with others’ ideas with medium engagement. This sequence demonstrates the capacity of the system to identify sophisticated interactions where students progressively build upon and critically engage with each other’s mathematical reasoning.
Table 4 demonstrates the system’s capacity to distinguish between mathematical discourse and off-topic conversations within classroom settings. While off-topic remarks are classified as uncorrelated, mathematical explanations receive relevant classifications such as engage with others’ ideas. The example further illustrates how the system recognizes varying levels of engagement with peer contributions, identifying both medium- and low-level engagement instances. Subsequently, when students transition to explaining their own reasoning, the system correctly adjusts the classification accordingly. This sequence exemplifies the system’s ability to maintain classification accuracy across diverse types of classroom discourse while maintaining focus on mathematically relevant interactions.
  Name Utterances Label Engagement Level
@@
Natalie I’m looking at the hundreds chart. (refers to yellow sheet) And then I went to 75. uncorrelated -
@@
Kimberly It was 70. uncorrelated -
Natalie On the second strategy. uncorrelated -
@@
Kimberly We are at the strategies, but she and me realized that we strategies on the 200 chart. explain own idea -
@@
2.4 Validation of Utterance Classification
To evaluate the reliability and validity of our utterance classification framework, we implemented a comprehensive multi-faceted validation approach. This validation strategy includes three primary analytical components: inter-model consistency analysis, human annotation comparison, and uncertainty quantification analysis. The inter-model consistency analysis assesses agreement levels between different LLMs, while human annotation comparison examines the consistency between human judgments and LLM-generated classifications. Additionally, we conducted both inter-model and intra-model uncertainty analyses to systematically identify instances requiring further review and to examine individual model confidence levels. These complementary approaches provide a thorough assessment of classification quality.
2.4.1 Inter-Model Consistency Analysis
We conducted a comprehensive inter-model consistency analysis to evaluate the degree of agreement between LLMs in their classification decisions. This analysis employed two established agreement metrics that provide complementary perspectives on model alignment: Cohen’s $\kappa$ (Cohen, 1960) and Fleiss’ $\kappa$ (Fleiss, 1971).
Cohen’s kappa
We utilized Cohen’s $\kappa$ coefficient (Cohen, 1960) to assess pairwise agreement between individual model pairs, which enables detailed examination of specific model relationships and identification of which models demonstrate the strongest concordance in their classification patterns.
Specifically, Cohen’s $\kappa$ coefficient $\kappa_{Cohen}$ quantifies the degree of agreement between two classification models while accounting for chance agreement. Cohen’s $\kappa$ coefficient is defined as:
$\kappa_{Cohen}=\frac{p_{o}-p_{e}}{1-p_{e}}$
where $p_{o}$ represents the observed proportional agreement between the two models under comparison, while $p_{e}$ denotes the expected proportional agreement that would occur due to chance alone. The metric provides a chance-corrected measure of agreement that operates within a standardized range from -1, indicating perfect disagreement, to +1, representing perfect agreement. A coefficient value of 0 means that the observed agreement is equivalent to what would be expected through random chance alone. The interpretation of Cohen’s $\kappa$ values follows established conventions as originally proposed by Landis and Koch (1977), with specific thresholds and their corresponding agreement levels detailed in Table 5.
$\kappa$ Statistic Strength of Agreement
<< 0.00 Poor
0.00 - 0.20 Slight
0.21 - 0.40 Fair
0.41 - 0.60 Moderate
0.61 - 0.80 Substantial
0.81 - 1.00 Almost perfect
Table 5: Standardized interpretation framework for Cohen’s $\kappa$ coefficient values, providing qualitative assessments of inter-rater agreement strength.
Figure 1 presents a heatmap visualization of pairwise Cohen’s $\kappa$ coefficients computed across all possible model combinations in our evaluation set. The analysis demonstrates predominantly strong agreement levels throughout the model pairs, with the majority of coefficients achieving substantial to almost perfect agreement ranges ($\kappa_{Cohen}>0.61$). However, a notable pattern emerges in the performance of Llama-3.1, which consistently exhibits lower agreement scores compared to all other models in the ensemble. This systematic deviation indicates that Llama-3.1 employs fundamentally different classification criteria or response generation mechanisms compared to the remaining six models.
Figure 1: Heatmap visualization displaying pairwise Cohen’s $\kappa$ coefficients computed across all possible combinations of the seven LLMs included in our ensemble framework. Each cell quantifies the chance-corrected agreement between two specific models, with color intensity representing the magnitude of agreement strength. Darker color intensities correspond to higher agreement levels, while lighter shades indicate lower consensus between model pairs.
Fleiss’ kappa
Additionally, we employed Fleiss’ $\kappa$ (Fleiss, 1971) to evaluate multi-rater agreement across all seven models simultaneously, which yields a measure of overall ensemble consistency.
Specifically, Fleiss’ $\kappa$ extends Cohen’s $\kappa$ framework to accommodate multiple raters in a single comprehensive analysis. Fleiss’ $\kappa$ is defined as:
$\kappa_{Fleiss}=\frac{\bar{P}-\bar{P_{e}}}{1-\bar{P_{e}}}$ (1)
where $\bar{P}$ represents the mean observed agreement across all classification items, while $\bar{P_{e}}$ denotes the mean expected agreement that would occur under the statistical assumption of independence among all seven models. This metric captures ensemble-wide consistency patterns that cannot be detected through pairwise comparisons alone, providing critical insights into the overall coherence.
Group Fleiss’ $\kappa$
LLMs (All) 0.6575
Open-source models 0.6033
Commercial API models 0.7354
Table 6: Fleiss’ $\kappa$ coefficients demonstrating multi-rater agreement levels across different model grouping. The comprehensive analysis includes three distinct categories: the complete seven-model ensemble represented by ”LLMs (All),” the subset of open-source models (“Open-source Models”), and the collection of commercial API models (“Commercial API Models”). All calculated coefficients achieve substantial agreement levels within the established range of 0.61-0.80, providing empirical evidence of reliable consensus across diverse model architectures and confirming the robustness of our classification methodology regardless of model type or commercial availability.
This comprehensive inter-model consistency analysis demonstrates robust agreement across our evaluation framework, with all models achieving substantial levels of consensus that validate the reliability of our classification approach. Fleiss’s kappa values consistently exceeded 0.60 across different model groupings, indicating strong agreement that surpasses conventional thresholds for acceptable inter-rater reliability. In particular, commercial API models demonstrated superior consensus ($\kappa_{Fleiss}$ = 0.7354) compared to their open source counterparts ($\kappa_{Fleiss}$ = 0.6033), suggesting that standardized development practices and potentially shared training methodologies within commercial ecosystems contribute to more aligned classification behaviors. The consistent identification of Llama-3.1 as an outlier in pairwise comparisons is a significant finding that merits further investigation of its underlying architectural or training characteristics that differentiate its classification approach from the majority of the ensemble. These results provide strong empirical support for the validity of our multi-model evaluation framework and establish confidence in the robustness and generalizability of our experimental findings across diverse model architectures and development paradigms.
2.4.2 Human-LLM Classification Comparison
To validate the reliability of our automated classification framework, we additionally applied a human annotation assessment and examined the consistency between human judgments and LLM-generated results. Specifically, we created standardized labeling guidelines that align precisely with the classification criteria used in our LLM prompts, ensuring methodological consistency between the automated and manual evaluation processes. Eight human annotators, comprising seven graduate students and one undergraduate student, participated in the validation study using a standardized web-based annotation platform specifically designed to minimize inter-annotator variability and maintain consistent labeling conditions.
The human annotation dataset comprised six strategically selected dialogue episodes that maximized the representation of the participants in classroom interactions. This focused subset was selected due to time and resource constraints, as comprehensive annotation of the complete dataset collected during the five-month instructional period would have exceeded the available capacity. To establish quantitative measures of annotation reliability, we implemented both Cohen’s $\kappa$ and Fleiss’ $\kappa$ coefficients for systematic assessment of inter-annotator agreement patterns. The heatmap visualization displaying pairwise Cohen’s $\kappa$ values between human annotators and LLMs is presented in Section LABEL:supp-sec:uncertainty of the Supplementary Material.
Group Fleiss’ $\kappa$
Human Annotators 0.3509
LLMs (All) 0.5070
Commercial API Models 0.5933
Table 7: Fleiss’ $\kappa$ Coefficients for human annotator agreement. Results are based on the same 5 selected blocks (out of 33 total blocks) used for human coding, while Table 6 uses all utterances across the full dataset.
Table 7 presents the Fleiss’ $\kappa$ coefficients across distinct annotator groups, revealing significant variations in classification consistency. The analysis demonstrates that human annotators achieved the lowest inter-rater agreement ($\kappa_{Fleiss}$ = 0.3509), indicating substantial disagreement in their classification decisions. In contrast, the complete LLM ensemble exhibited moderate consistency levels ($\kappa_{Fleiss}$ = 0.5070), while commercial API models demonstrated the highest agreement ($\kappa_{Fleiss}$ = 0.5933) among all the evaluated groups. These findings illuminate the inherent challenges associated with dialogue classification tasks, where subjective interpretation and contextual ambiguity contribute to variability in human judgment. The observed pattern suggests that multiple valid interpretations may exist for individual utterances, reflecting the complex nature of conversational analysis rather than deficiencies in human annotation capabilities. The higher consistency demonstrated by automated classification systems, particularly commercial models, indicates that standardized algorithmic approaches may provide more reproducible labeling results compared to human annotators.
2.4.3 Uncertainty Quantification
Inter-Model Uncertainty Quantification
To systematically quantify the degree of disagreement among models and identify utterances requiring expert review, we implemented Shannon entropy (Shannon, 1948) as our primary uncertainty measure. This information-theoretic approach enables consistent measurement of classification uncertainty by analyzing the probability distribution of label choices across multiple LLM outputs, where each label’s probability is calculated as the proportion of models selecting that label. For a given utterance with classification probabilities $\{p_{1},p_{2},\dots,p_{n}\}$ distributed across $n$ possible classification labels, the Shannon entropy, $H(X)$, is defined as:
$H(X)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i},$ (2)
where the probability $p_{i}$ for each label category is empirically estimated through ensemble voting as: